A bison is traveling at 19m/s when it collides with an elephant going 12m/S at 40°. What is the impact velocity relative to the bison


1
Expert's answer
2021-09-14T09:35:51-0400
"v=\sqrt{19^2+12^2-2(19)(12)\cos{40}}=12.5\frac{m}{s}"